OpenStudy (anonymous):

Simplify: 14-2[9-2(5-3)]

OpenStudy (mathstudent55):

Start working on the innermost parentheses: 5 - 3 = 2

OpenStudy (anonymous):

work first on the innermost parenthesis which is 5 - 3

OpenStudy (anonymous):

then multiply it by 2

OpenStudy (mathstudent55):

14-2[9-2(5-3)] = 14-2[9-2(2)] =

OpenStudy (anonymous):

then move on to the next parenthesis 9 - 4

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

then multiply by 2 u get 8 then subtract 14 and 8 u get 6 :)

OpenStudy (mathstudent55):

14-2[9-2(5-3)] = 14-2[9-2(2)] = 14 -2[9 - 4] = 14 - 2[5] = 14 - 10 = 4

OpenStudy (mathstudent55):

@delarican95 Are you familiar with PEMDAS, the order of operations?